Two ships leave one port, one sailing east and the other south. Some time later, they are 17 mies apart, with the eastbound ship 7 miles farther from the port than the southbound ship. How far is each from port?
Answer by mananth(16946)

let southbound ship be at x miles from port
the other will be (x+7) miles from the port east.
they form a right triangle
x^2+(x+7)^2=17^2
x^2+x^2+14x+49=289
2x^2+14x=289-49
2x^2+14x=240
divide by 2
x^2+7x=120
x^2+7x-120=0
x^2+15x-8x-120=0
x(x+15)-8(x+15)=0
(x+15)(x-8)=0
x= -15 or x = 8 ignore negative value.
x=8 miles south bound ship from port
x+7 = 8+7 =15 miles distance of the east bound ship from port